OpenStudy (anonymous):

helpppp...... Complete this item. (Enter letter variables in alphabetical order.) write 3a^2b^3c^5 / 8x^4y^3z using no denominator. (No fraction line is necessary. use * for multiplication between numbers.)

OpenStudy (phi):

use the rule \[ \frac{1}{a^n} = a^{-n} \] can you give it a try ?

OpenStudy (anonymous):

i don't know were i should start.....

OpenStudy (phi):

how about with the 8 ? \[ \frac{1}{8} \text{ is the same as } \frac{1}{8^1} \] use the rule \[ \frac{1}{a^n} = a^{-n} \] what do you get ?

OpenStudy (phi):

Do you know how to interpret the rule ? "pattern match" your problem with the rule notice that the 8 matches with the "a" and the little 1 in the upper right of 8 matches with the "n"

OpenStudy (anonymous):

sorry don't know what u r saying....

OpenStudy (anonymous):

whats the answer?????